# An Interpolation of Metrics and Spaces of Metrics

### v2

@@ 1.1. Backgrounds
The function $\mathcal{D}_{X}$ is a metric on $\mathrm{Met}(X)$ valued in $[0,\infty]$. Throughout this paper, we consider that $\mathrm{Met}(X)$ is always equipped with the topology generated by $\mathcal{D}_{X}$. To investigate topological properties of $\mathrm{Met}(X)$, we first generalize the Hausdorff extension theorem to an interpolation theorem of metrics with an approximation by $\mathcal{D}_{X}$ (Theorem 1.1 or 3.4). As applications of it, for a certain property $\mathcal{P}$ on metric spaces, we prove that the set set $\{\,d\in\mathrm{Met}(X)\mid\text{$(X,d)$ satisfies $\mathcal{P}$}\,\}$ is generic in $\mathrm{Met}(X)$ (see Theorems 1.2 and 1.3). This result can be considered as an analogue of Banach’s famous result, which states that in the continuous function space on $[0,1]$, a generic function is nowhere differentiable.
@@ 1.2.1. Interpolation of metrics
For every metrizable space $X$, and for every closed subset $A$ of $X$, Nguyen Van Khue and Nguyen To Nhu [16] constructed a Lipschitz metric extensor from $(\mathrm{Met}(A),\mathcal{D}_{A})$ into $(\mathrm{Met}(X),\mathcal{D}_{X})$, and a monotone continuous metric extensor from $\mathrm{Met}(A)$ into $\mathrm{Met}(X)$; moreover, if $X$ is completely metrizable, then each of these metric extensors maps any complete metric in $\mathrm{Met}(A)$ into a complete metric in $\mathrm{Met}(X)$. To obtain such metric extensors, they used the Dugundji extension theorem concerning locally convex topological linear spaces.
@@ 2.3. The Kuratowski embedding theorem
For a metric space $(X,d)$, we denote by $C_{b}(X)$ the Banach space of all bounded continuous functions on $X$ equipped with the supremum norm. For $x\in X$, we denote by $d_{x}$ the function from $X$ to $\mathbb{R}$ defined by $d_{x}(p)=d(x,p)$. The following theorem, which states that every metric space is isometrically embeddable into Banach spaces, is known as the Kuratowski embedding theorem. The proof can be seen in, for example, [8, Theorem 4.3.14].
@@ 2.4. Baire spaces
Since $G_{\delta}$ subset of completely metrizable space is completely metrizable (see, e.g. [23, Theorem 24.12]), we obtain the following:
@@ 3.1. Amalgamation lemmas
Then $h$ is a metric as desired. For more details, we refer the readers to [5, Theorem 7.3.25]. ∎
@@ 4. Transmissible properties
$d_{L}(x,y)=\begin{cases}d_{i}(x,y)&\text{if $x,y\in R_{i}$ for some $i$;}\\ \epsilon\cdot\max\{2^{-i},2^{-j}\}&\text{if $x\in R_{i},y\in R_{j}$ for some $i\neq j$; }\\ \epsilon\cdot 2^{-i}&\text{if $x=\infty,y\in R_{i}$ for some $i$;}\\ \epsilon\cdot 2^{-i}&\text{if $x\in R_{i},y=\infty$ for some $i$.}\end{cases}$
On account of (R1), $d_{L}$ is actually a metric (compare with [14, Definition 3.3]). We also observe that the space $(L,d_{L})$ is an $(\omega_{0}+1)$-metric space with $\delta_{d_{L}}(L)\leq\epsilon$. By the properties (R2) and (R3) of $\{(R_{i},d_{i})\}_{i\in\mathbb{N}}$, the metric space $(L,d_{L})$ satisfies the anti-$\mathfrak{G}$-transmissible property. ∎
@@ 4.2. The doubling property and the uniform disconnectedness
Note that $(X,d)$ is doubling if and only if $(X,d)$ has finite Assouad dimension (see e.g., [13, Section 10]).
@@
A metric space $(X,d)$ is said to be uniformly disconnected if there exists $\delta\in(0,1)$ such that if a finite sequence $\{z_{i}\}_{i=1}^{N}$ in $X$ satisfies $\max_{i}d(z_{i},z_{i+1})<\delta d(z_{1},z_{N})$, then we have $N=1$. Namely, the inequality $N>1$ implies $\max_{i}d(z_{i},z_{i+1})\geq\delta d(z_{1},z_{N})$. Note that a metric space is uniformly disconnected if and only if it is bi-Lipschitz to an ultrametric space (see e.g., [18, Lemma 5.1.10]).
@@ 4.3. Rich pseudo-cones
Let $(X,d)$ be a metric space. Let $\{A_{i}\}_{i\in\mathbb{N}}$ be a sequence of subsets of $X$, and let $\{u_{i}\}_{i\in\mathbb{N}}$ be a sequence in $(0,\infty)$. We say that a metric space $(P,d_{P})$ is a pseudo-cone of $X$ approximated by ($(\{A_{i}\}_{i\in\mathbb{N}},\{u_{i}\}_{i\in\mathbb{N}})$ if
@@
For a metric space $(X,d)$, the $-\mathfrak{R}$-transmissible property means that for every $n\in\mathbb{N}$, and for every $m\in\mathbb{N}$, there exist a finite subspace $A=\{a_{i}\}_{i=1}^{\operatorname{card}(L_{n})}$ of $X$ and a positive number $z\in(0,\infty)$ such that for all $i,j\in\{1,\dots,\operatorname{card}(L_{n})\}$ we have
@@
in particular, $\mathcal{GH}((A,z^{-1}d|_{A^{2}}),(L_{n},d_{n}))<2^{-(m+1)}$. Thus $\mathscr{F}$ is contained in $\operatorname{\mathscr{P}}(X,d)$. This implies that $(X,d)$ has rich pseudo-cones. We next prove the opposite. If $(X,d)$ has rich pseudo-cones, then for every $(W,d_{W})\in\mathscr{F}$, and for every $\epsilon\in(0,\infty)$, there exist a positive number $z\in(0,\infty)$ and a subset $A$ of $X$ with $\operatorname{card}(A)=\operatorname{card}(W)$ such that $\mathcal{GH}((A,z^{-1}d|_{A^{2}}),(W,d_{W}))<\epsilon$. Thus $(X,d)$ satisfies the $-\mathfrak{R}$-transmissible property. We next prove that $\mathfrak{R}$ is singular. For each $(n,m)\in\mathbb{N}^{2}$ and for each $\epsilon\in(0,\infty)$, we put
@@ 4.4. Metric inequality
    where the symbol $+^{\prime\prime}$ stands for the addition of $\mathbb{Z}/m\mathbb{Z}$.
@@ 5. Local Transmissible properties
Using the assumption that $X$ is second-countable and locally compact, now we take a sequence $\{B(n)\}_{n\in\mathbb{N}}$ of compact subsets of $X^{2}$ with $\bigcup_{n\in\mathbb{N}}B(n)=X^{2}\setminus\Delta_{X}$, where $\Delta_{X}$ is the diagonal set of $X^{2}$, and take a sequence $\{K(n)\}_{n\in\mathbb{N}}$ of compact subsets of $X$ with $K(n)\subset\operatorname{INT}(K(n+1))$ and $\bigcup_{n\in\mathbb{N}}K(n)=X$, where “$\operatorname{INT}$” means the interior.
@@
00 0 $<d(K(n),X\setminus\operatorname{INT}(K(n+1)))\leq d(K(n),X\setminus K(n+1))$

### v3

@@ 1.1. Backgrounds
The function $\mathcal{D}_{X}$ is a metric on $\mathrm{Met}(X)$ valued in $0,\infty]$. Throughout this paper, we consider that $\mathrm{Met}(X)$ is always equipped with the topology generated by $\mathcal{D}_{X}$. To investigate topological properties of $\mathrm{Met}(X)$, we first generalize the Hausdorff extension theorem to an interpolation theorem of metrics with an approximation by $\mathcal{D}_{X}$ (Theorem 1.1 or 3.4). As applications of it, for a certain property $\mathcal{P}$ on metric spaces, we prove that the set set $\{\,d\in\mathrm{Met}(X)\mid\text{$(X,d)$ satisfies $\mathcal{P}$}\,\}$ is generic in $\mathrm{Met}(X)$ (see Theorems 1.2 and 1.3). This result can be considered as an analogue of Banach’s famous result, which states that in the continuous function space on $[0,1]$, a generic function is nowhere differentiable.
@@ 1.2.1. Interpolation of metrics
For every metrizable space $X$, and for every closed subset $A$ of $X$, Nguyen Van Khue and Nguyen To Nhu [20] constructed a Lipschitz metric extensor from $(\mathrm{Met}(A),\mathcal{D}_{A})$ into $(\mathrm{Met}(X),\mathcal{D}_{X})$, and a monotone continuous metric extensor from $\mathrm{Met}(A)$ into $\mathrm{Met}(X)$; moreover, if $X$ is completely metrizable, then each of these metric extensors maps any complete metric in $\mathrm{Met}(A)$ into a complete metric in $\mathrm{Met}(X)$. To obtain such metric extensors, they used the Dugundji extension theorem concerning locally convex topological linear spaces. For more information of simultaneous extensions of metrics, see [3], [5], [4], [29], [25], [26], [24], and [17]. It seems difficult to prove Theorem 1.1 using these extension result because they do not preserve a given metric $d$.
@@ 2.3. The Kuratowski embedding theorem
For a metric space $(X,d)$, we denote by $C_{b}(X)$ the Banach space of all bounded continuous functions on $X$ equipped with the supremum norm. For $x\in X$, we denote by $d_{x}$ the function from $X$ to $\mathbb{R}$ defined by $d_{x}(p)=d(x,p)$. The following theorem, which states that every metric space is isometrically embeddable into Banach spaces, is known as the Kuratowski embedding theorem. The proof can be seen in, for example, [11, Theorem 4.3.14].
@@ 2.4. Baire spaces
Since $G_{\delta}$ subset of completely metrizable space is completely metrizable (see, e.g. [31, Theorem 24.12]), we obtain the following:
@@ 3.1. Amalgamation lemmas
Then $h$ is a metric as desired. For more details, we refer the readers to [8, Theorem 7.3.25]. ∎
@@ 4. Transmissible properties
[…]
On account of (R1), $d_{L}$ is actually a metric (compare with [18, Definition 3.3]). We also observe that the space $(L,d_{L})$ is an $(\omega_{0}+1)$-metric space with $\delta_{d_{L}}(L)\leq\epsilon$. By the properties (R2) and (R3) of $\{(R_{i},d_{i})\}_{i\in\mathbb{N}}$, the metric space $(L,d_{L})$ satisfies the anti-$\mathfrak{G}$-transmissible property. ∎
@@ 4.2. The doubling property and the uniform disconnectedness
Note that $(X,d)$ is doubling if and only if $(X,d)$ has finite Assouad dimension (see e.g., [16, Section 10]).
@@
A metric space $(X,d)$ is said to be uniformly disconnected if there exists $\delta\in(0,1)$ such that if a finite sequence $\{z_{i}\}_{i=1}^{N}$ in $X$ satisfies $\max_{i}d(z_{i},z_{i+1})<\delta d(z_{1},z_{N})$, then we have $N=1$. Namely, the inequality $N>1$ implies $\max_{i}d(z_{i},z_{i+1})\geq\delta d(z_{1},z_{N})$. Note that a metric space is uniformly disconnected if and only if it is bi-Lipschitz to an ultrametric space (see e.g., [22, Lemma 5.1.10]).
@@ 4.3. Rich pseudo-cones
Let $(X,d)$ be a metric space. Let $\{A_{i}\}_{i\in\mathbb{N}}$ be a sequence of subsets of $X$, and let $\{u_{i}\}_{i\in\mathbb{N}}$ be a sequence in $(0,\infty)$. We say that a metric space $(P,d_{P})$ is a pseudo-cone of $X$ approximated by $(\{A_{i}\}_{i\in\mathbb{N}},\{u_{i}\}_{i\in\mathbb{N}})$ if
@@
For a metric space $(X,d)$, the anti-$\mathfrak{R}$-transmissible property means that for every $n\in\mathbb{N}$, and for every $m\in\mathbb{N}$, there exist a finite subspace $A=\{a_{i}\}_{i=1}^{\operatorname{card}(L_{n})}$ of $X$ and a positive number $z\in(0,\infty)$ such that for all $i,j\in\{1,\dots,\operatorname{card}(L_{n})\}$ we have
@@
in particular, $\mathcal{GH}((A,z^{-1}d|_{A^{2}}),(L_{n},d_{n}))<2^{-(m+1)}$. Thus $\mathscr{F}$ is contained in $\operatorname{\mathscr{P}}(X,d)$. This implies that $(X,d)$ has rich pseudo-cones. We next prove the opposite. If $(X,d)$ has rich pseudo-cones, then for every $(W,d_{W})\in\mathscr{F}$, and for every $\epsilon\in(0,\infty)$, there exist a positive number $z\in(0,\infty)$ and a subset $A$ of $X$ with $\operatorname{card}(A)=\operatorname{card}(W)$ such that $\mathcal{GH}((A,z^{-1}d|_{A^{2}}),(W,d_{W}))<\epsilon$. Thus $(X,d)$ satisfies the anti-$\mathfrak{R}$-transmissible property. We next prove that $\mathfrak{R}$ is singular. For each $(n,m)\in\mathbb{N}^{2}$ and for each $\epsilon\in(0,\infty)$, we put
@@ 4.4. Metric inequality
    where the symbol $^{\prime\prime}$ stands for the addition of $\mathbb{Z}/m\mathbb{Z}$.
@@ 5. Local Transmissible properties
Using the assumption that $X$ is second-countable and locally compact, now we take a sequence $\{B(n)\}_{n\in\mathbb{N}}$ of compact subsets of $X^{2}$ with $\bigcup_{n\in\mathbb{N}}B(n)=X^{2}\setminus\Delta_{X}$, where $\Delta_{X}$ is the diagonal set of $X^{2}$, and take a sequence $\{K(n)\}_{n\in\mathbb{N}}$ of compact subsets of $X$ with $K(n)\subset\operatorname{INT}(K(n+1))$ and $\bigcup_{n\in\mathbb{N}}K(n)=X$, where $\operatorname{INT}$” means the interior.
@@
$0$ $<d(K(n),X\setminus\operatorname{INT}(K(n+1)))\leq d(K(n),X\setminus K(n+1))$